# Sparsity May Be All You Need: Sparse Random Parameter Adaptation

### v2

@@ Appendix E Training Hyper-Parameters
The hyper-parameters used for investigating SpaRTA and other adaptation methods on IMDB and GLUE are summarized next.
@@ E.1 IMDB
Model: google/gemma-2b-it method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-4 5e-5 5e-5 1e-5 2e-4 5e-5 5e-5 SpaRTA (0.5%) 2e-4 1e-3 2e-4 2e-4 2e-4 5e-4 1e-3 2e-4 SpaRTA (0.037%) 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 LoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 DoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4
@@
Model: Mistral-7B-Instruct-v0.3 method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-6 5e-6 1e-5 1e-5 5e-5 5e-5 5e-6 SpaRTA (0.5%) 2e-4 1e-4 5e-5 5e-4 2e-4 5e-4 2e-4 5e-5 SpaRTA (0.048%) 5e-4 5e-4 5e-4 1e-3 5e-4 1e-3 1e-3 1e-3 LoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 DoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4
@@
Table 16: (IMDB) Training parameters used with each fine-tuning method and model in Table 4. $\ast$ in number of epochs indicates early stopping was used. For SpaRTA, parameters for density 5%, 0.5% 0,05% are reported.
@@ E.2 GLUE and BoolQ
For GLUE and BoolQ, training was done using a simple grid search over the learning rates: [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6]. The optimal learning rates found are given in Tables 14 and 15. Weight decay was set to 0. Dropout was set to 0.0 for SpaRTA and 0.1 for LoRA and DoRA. Batch sizes were set according to what could fit in GPU memory (different per dataset and model). The number of epochs was set to 3 for instruct models and 2 for base models. Similarly to IMDB, any sample longer than 256 tokens (which is dependnent on the tokenizer used) is discarded from the dataset to avoid having few minibatch with one very long sample compared to the rest, as seen in Table 2.

### v3

@@ Appendix E Training Hyper-Parameters
The hyper-parameters used for investigating SpaRTA and other adaptation methods are summarized next.
@@ E.1 IMDB
Model: gemma-2b-it Method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-4 5e-5 5e-5 1e-5 2e-4 5e-5 5e-5 SpaRTA (0.5%) 2e-4 1e-3 2e-4 2e-4 2e-4 5e-4 1e-3 2e-4 SpaRTA (0.037%) 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 LoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 DoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4
@@
Model: mistral-7b-it Method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-6 5e-6 1e-5 1e-5 5e-5 5e-5 5e-6 SpaRTA (0.5%) 2e-4 1e-4 5e-5 5e-4 2e-4 5e-4 2e-4 5e-5 SpaRTA (0.048%) 5e-4 5e-4 5e-4 1e-3 5e-4 1e-3 1e-3 1e-3 LoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 DoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4
@@
Table 19: (IMDB) Training parameters used with each fine-tuning method and model in Table 4. An $\ast$ in number of epochs indicates early stopping was used. For SpaRTA, parameters for density 5%, 0.5% 0,05% are reported.
@@ E.2 GLUE and BoolQ
For GLUE and BoolQ, training was done using a simple grid search over the learning rates: [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6]. The optimal learning rates found are given in Tables 17 and 18. Weight decay was set to 0. Dropout was set to 0.0 for SpaRTA and 0.1 for LoRA and DoRA. Batch sizes were set according to what could fit in GPU memory (different per dataset and model). The number of epochs was set to 3 for instruct models and 2 for base models. Similarly to IMDB, any sample longer than 256 tokens (which is dependnent on the tokenizer used) is discarded from the dataset to avoid having few minibatch with one very long sample compared to the rest, as seen in Table 2.